The rectangle has a perimeter of 78 centimeters.the length of the rectangle is 11 centimeters longer than triple the width. Find the demensions of the rectangle
Answer by rapaljer(4671)

Let x = width
3x +11 = length

2(x) + 2(3x+11) = 78
2x + 6x + 22 = 78
8x + 22 = 78
8x + 22 - 22 = 78 - 22
8x = 56
x= 7 Width
3x+11 = 3(7) + 11 = 32 Length

Check:
2(7) + 2(32)= 78
14 + 64 = 78
It checks!

R^2 at SCC